OpenStudy (anonymous):

88 children were in a swimming camp. 1/3 of the boys and 3/7 of the girls wore goggles. If 34 students wore goggles, how many girls wore goggles?

OpenStudy (anonymous):

Let the n.o of boys be x and girls be y 2/3x+4/7y=54 1/3x+3/7y=34 1/3x=34-3/7y Sub this in 2(34-3/7y)+4/7y=54 you solve it and get y=49 n.o of girls who wore goggles=49 x 3/7=21